OpenStudy (anonymous):

is 2x^2-x^2 equal to 2

OpenStudy (anonymous):

\[2x ^{2}-x ^{2}=x ^{2}\]

OpenStudy (anonymous):

Thank-you

geerky42 (geerky42):

Here's the tip; notice that both terms have x² so you can factor it out: \[\Large 2x^2-x^2 = x^2(2-1) = x^2(1) = \boxed{x^2}\]